You’re riding your bicycle first west a distance of 10.6 km, then south a distance of 14.1 km.

(a)

In what direction does your average velocity vector point, in degrees south of west?

(b)

If the total trip took 10.3 minutes, what is the magnitude of this average velocity, in m/s?
Answer #1

a)

direction

x = arctan ( 14.1 / 10.6)

x = 53.07

=======

b)

net displacement

d^2 = 14.1^2 + 10.1^2

d = 17.344 m

av velocity = d /t = 17.344 / ( 10.3* 60)

v = 0.028 m/s

======

Comment before rate in case any doubt, will reply for sure.. goodluck